OpenStudy (anonymous):

Find the area of the region between a regular hexagon with sides of 6" and its inscribed circle

OpenStudy (anonymous):

I think I have the answer, but Im not sure. I got \[10\sqrt{3}-27\pi \] square inches

OpenStudy (akashdeepdeb):

Is that 6 inches?

OpenStudy (anonymous):

are you asking if it says 6" in the question?

OpenStudy (anonymous):

If so, then yes.

OpenStudy (akashdeepdeb):

Yes! :D

OpenStudy (akashdeepdeb):

Okay! Let me just see! :D

OpenStudy (anonymous):

the answer needs to be formatted \[anumber \times \sqrt{anumber}-?\times \pi\]

OpenStudy (akashdeepdeb):

 Did you get the answer like this? :D

OpenStudy (anonymous):

yes!

OpenStudy (akashdeepdeb):

Then it should be CORRECT!! :D

OpenStudy (anonymous):

thank you so much!

OpenStudy (akashdeepdeb):

:)